# Spin Photonic Topological Insulator Leaky Wave Antenna By Armchair Arrangement

### v3

@@
A new X-band leaky wave antenna has been developed using spin photonic topological insulators. This antenna features a hexagonal unit cell arranged in an armchair configuration. This arrangement provides advantages over the zigzag configuration, particularly by offering a wider leaky region and a more suitable pattern. To design the structure, a parametric study on the cell dimensions has been conducted, and another study has designed the transition region to couple the topological structure with the classical line. The proposed antenna has a low profile and illuminates two sides of the structure simultaneously. Additionally, it offers a 53-degree scanning range and a bandwidth of 2.7 GHz, making it a groundbreaking improvement over other leaky wave antennas that utilize photonic topological insulators. The proposed antenna is a unbalaned CRLH leaky wave antenna that can radiate in both backward and forward directions. Notably, when transitioning through the broadside in the scanning range, there is no significant drop in performance despite the existence of an open stop band. To the best of our knowledge, this characteristic appears to be unique among unbalaned CRLH leaky wave antennas.
@@ 1 Introduction
This paper proposes that adopting an armchair arrangement, instead of a zigzag design, effectively addresses the identified issues and facilitates backward-to-forward beam scanning. A crucial point is that broadside radiation in unbalanced composite right/left-handed (CRLH) leaky wave antennas presents significant challenges, often leading to a sudden drop in gain due to an open stopband. However, in the proposed antenna with the armchair arrangement, the impact of this phenomenon is significantly reduced. Simulations were conducted using Ansys HFSS and were subsequently repeated in CST Studio to validate the results.
@@ 3 Antenna application
The main advantage of using the armchair arrangement in antenna applications is that it effectively places most of the topological bandgap within the fast wave region. According to the findings in Appendix A, this design enables the proposed structure to scan through the broadside, a capability that conventional unbalaned CRLH leaky wave antennas do not offer.
@@ 2 The broadside is at $90^{\circ}$
At 15.8 GHz, the realized gain in the open stopband is less than -3 dB compared to the values observed at the ends of the scanning range. To the best of our knowledge, this unusual phenomenon does not occur in other unbalaned CRLH leaky wave antennas. The 36-degree scan at both the top and bottom of the antenna surpasses the findings reported in [33], matching those in [32], even as it crosses the broadside.

### v4

@@
A new X-band leaky wave antenna has been developed using spin photonic topological insulators. This antenna features a hexagonal unit cell arranged in an armchair configuration. This arrangement provides advantages over the zigzag configuration, particularly by offering a wider leaky region and a more suitable pattern. To design the structure, a parametric study on the cell dimensions has been conducted, and another study has designed the transition region to couple the topological structure with the classical line. The proposed antenna has a low profile and illuminates two sides of the structure simultaneously. Additionally, it offers a 53-degree scanning range and a bandwidth of 2.7 GHz, making it a groundbreaking improvement over other leaky wave antennas that utilize photonic topological insulators. The proposed antenna is an unbalanced Composite Right/Left-Handed (CRLH) leaky wave antenna capable of radiating in both backward and forward directions. Notably, as it transitions through the broadside within its scanning range, there is no significant drop in performance, even in the presence of an open stop band. To the best of our knowledge, this characteristic is unique among unbalanced CRLH leaky wave antennas.
@@ 1 Introduction
This paper proposes that adopting an armchair arrangement, instead of a zigzag design, effectively addresses the identified issues and facilitates backward-to-forward beam scanning. A crucial point is that broadside radiation in unbalanced CRLH leaky wave antennas presents significant challenges, often leading to a sudden drop in gain due to an open stopband. However, in the proposed antenna with the armchair arrangement, the impact of this phenomenon is significantly reduced. Simulations were conducted using Ansys HFSS and were subsequently repeated in CST Studio to validate the results.
@@ 3 Antenna application
The main advantage of using the armchair arrangement in antenna applications is that it effectively places most of the topological bandgap within the fast wave region. According to the findings in Appendix A, this design enables the proposed structure to scan through the broadside, a capability that conventional unbalanced CRLH leaky wave antennas do not offer.
@@ 2 The broadside is at $90^{\circ}$
At 15.8 GHz, the realized gain in the open stopband is less than -3 dB compared to the values observed at the ends of the scanning range. To the best of our knowledge, this unusual phenomenon does not occur in other unbalanced CRLH leaky wave antennas. The 36-degree scan at both the top and bottom of the antenna surpasses the findings reported in [33], matching those in [32], even as it crosses the broadside.